OpenStudy (anonymous):

At maximum speed, an airplane travels 1680 miles against the wind in 5 hours. Flying with the wind, the plane can travel the same distance in 4 hours. What is the speed of the plane with no wind?

OpenStudy (anonymous):

(miles / hour) * (hours) = miles miles/hours = miles per hour

OpenStudy (anonymous):

Call speed of plane x, speed of wind y distance=speed*time first equation for against the wind 1680=(x-y)5 second for with the wind 1680=(x+y)4 now solve as a system of equations to find x

OpenStudy (anonymous):

Thank you!!!!